I need to work out a ratio problem


Linda and Blake collected money for charity in the ratio of 5:2 respectively. Blake collected 600 less than Linda. Determine the total amount collected by Linda and Blake.

Please could you explain how I go about working out this problem


1
Expert's answer
2021-03-11T06:17:27-0500

Linda and Blake collected money in the ratio = 5:2

We can conclude from the ratio that Linda collected money as 5x and Blake collected money as 2x.


Given that Blake collected money 600 less than Linda but through the ratio we can conclude that Blake had collected money less than Linda is as : 5x - 2x = 3x

Thus, 3x = 600

or x = 200

The total amount collected by both Linda and Blake is = 5x + 2x = 7x

= 7*(200) = 1400